OpenStudy (anonymous):

Subtract: (8-3i)-(5-4i)

OpenStudy (anonymous):

8-3j-5+4j = 3+j

OpenStudy (anonymous):

8-5-3i+4i =3+i

OpenStudy (anonymous):

(8-3i)-(5-4i) (8-3i-5+4i) (8-5-3i+4i) (3+i) answer=3+i